OpenStudy (amenah8):

L-Hopital's Rule: lim x--> infinity of (e^3x)/(x^2)

jimthompson5910 (jim_thompson5910):

what happens when x approaches infinity? what form does (e^3x)/(x^2) take on?

OpenStudy (amenah8):

it is indeterminate: infinity/infinity

jimthompson5910 (jim_thompson5910):

correct, so `L-Hopital's Rule` does apply here

jimthompson5910 (jim_thompson5910):

I'm assuming the numerator is \(\LARGE e^{3x}\) right?

OpenStudy (amenah8):

yes

jimthompson5910 (jim_thompson5910):

what is the derivative of e^(3x) equal to?

OpenStudy (amenah8):

e^3x

jimthompson5910 (jim_thompson5910):

you forgot to apply the chain rule

jimthompson5910 (jim_thompson5910):

you need to derive the inner stuff (3x) to get 3 so if y = e^(3x) then dy/dx = 3*e^(3x)

OpenStudy (amenah8):

oh right!

jimthompson5910 (jim_thompson5910):

now take the derivative of the denominator x^2 to get ______

OpenStudy (amenah8):

2x

OpenStudy (amenah8):

then e^infinity / infinity)

jimthompson5910 (jim_thompson5910):

so we go from \[\Large \lim_{x \to \infty} \left[\frac{e^{3x}}{x^2}\right]\] to \[\Large \lim_{x \to \infty} \left[\frac{3e^{3x}}{2x}\right]\]

jimthompson5910 (jim_thompson5910):

if that second limit leads to infinity/infinity, you'll need to apply `L' Hospital's Rule` again

OpenStudy (amenah8):

so: e^3x (0) / 2

OpenStudy (amenah8):

as the derivative again?

jimthompson5910 (jim_thompson5910):

you should go from \[\Large \lim_{x \to \infty} \left[\frac{3e^{3x}}{2x}\right]\] to \[\Large \lim_{x \to \infty} \left[\frac{9e^{3x}}{2}\right]\]

jimthompson5910 (jim_thompson5910):

I don't think you applied the chain rule properly

OpenStudy (amenah8):

oh you're right! i see my mistake

OpenStudy (amenah8):

you get infinity / 2

OpenStudy (amenah8):

so the answer is infinity?

jimthompson5910 (jim_thompson5910):

anyways at this point, you should find that the entire expression now approaches infinity as x approaches infinity ie, \[\Large \lim_{x \to \infty} \left[\frac{9e^{3x}}{2}\right] = \infty\]

jimthompson5910 (jim_thompson5910):

yep, \[\Large \lim_{x \to \infty} \left[\frac{e^{3x}}{x^2}\right] = \infty\]

OpenStudy (amenah8):

thank you so much! this really helped!!

jimthompson5910 (jim_thompson5910):

no problem